please help me solve each system of equation
1.) y=3x-4 2.) x-2y=-14 3.) x+2y=4 4.) 3x-y=2
y=-2x+1 y+2x=4 2x+y=-1 y=2x-5
and whats the solution??

Answer by jeselle(2)

by substitution:
1.)
y=3x-4
y=-2x+1
simply substitute the value of y in the second equation to the first
-2x+1=3x-4
-2x-3x=-4-1
-5x=-5
(divide both sides by -5)
x=1
by elimination
-3x+y=-4
2x+y=1
(subtract so that the sign of the subtrahend will be the opposite)
-3x+y=-4
-1(2x+y=1)-1

-3x+y=-4
-2x-y=-1
then add
-5x=-5
x=1
do the same with the others and try practicing.^^